Consider a binomial experiment with

n = 20

and

p = 0.80.

(Round your answers to four decimal places.)

(a)

Compute

f(12).

       

f(12) =    

       

(b)

Compute

f(16).

       

f(16) =    

       

(c)

Compute

P(x ≥ 16).

       

P(x ≥ 16) =    

       

(d)

Compute

P(x ≤ 15).

       

P(x ≤ 15) =    

       

(e)

Compute

E(x).

       

E(x) =    

       

(f)

Compute

Var(x)

and σ.

                       

Var(x)

            =                                                       σ             =            
Answer #1

و ع دد = m م2 2 09-0) - (۵۰۶ (۱۰۰) . = (2 روعه 218e .1 = 4 (5 . 2) (2 006 ، ملط - - و 1، s1 )مل -20 (.) *0 . 1) 2 - ا= /90 2
Answer #2

Given ! 14 Xe gio omial (n = Le, P = 0. 7 0) f(12) = (38) (0.80112 (0.2008 +(12) = 0.0222 f(16) = (0(0.80)16 (0.2014 = 0.2182